Show your solution.

  A 12-volt battery will increase the potential energy of ____ coulombs of charge by 6 joules.


1
Expert's answer
2021-02-28T07:30:03-0500

Let's write the relationship between the electrical potential difference, potential energy and the amount of charge:


"\Delta V=\dfrac{W}{q}=\dfrac{\Delta PE}{q},""q=\dfrac{\Delta PE}{\Delta V}=\dfrac{6\ J}{12\ V}=0.5\ C."

Answer:

A 12-volt battery will increase the potential energy of 0.5 coulombs of charge by 6 joules.